Two concentric circular coils A and B
have radii 25cm and 15cm and carry currents 10A and 15A respoectively.A has 24 turns and B has 18 turns.the direction of the currents are in opposite order . the magnetic induction at the common center of the coil is ?

Grade:12

1 Answers

Pradeep Rawat
5 years ago
 

 for coil A: 

VA=25cm=0.25m

nA=24nA=24

IA=10AI=10A

For coil B:

rb=15cm=0.15m

nb=18nb=18

Ib=15AI=15A

Magnetic field at the center of coil x

BA=µ0 /4π*2Ia πna/ra

=10−7×2×10×3.14×24/0.25

=6.028×10^−4

=6.028×10^−4T

Magnetic field at the center of coil B

B= µ0 /4π*2Ib πnb/rb

=10−7×2×3.14×15×18

=11.30×10−4T

Hence the net magnetic field at the center

Bb−Ba=(11.30×10−4−6.028×10−4)T

=5.27×10−4T